You are applying for jobs at two companies. Company C offers starting salaries with μ = $39,000 and σ = $4000. Company D offers starting salaries with μ = $39,000 and σ = $1500. From which company are you more likely to get an offer of $42,000 or more? Explain your reasoning.
Answer by rothauserc(4718)

We use a normal distribution and calculate the z-score for each company
:
z-score for Company C = (42000 - 39000) / 4000 = 0.75
:
z-score for Company D = (42000 - 39000) / 1500 = 2.00
:
We consult z-tables for Probability(Pr)
:
Pr ( X > 42000 ) for C = 1 - Pr ( X < 42000 ) = 1 - 0.7734 = 0.2266
:
Pr ( X > 42000 ) for D = 1 - Pr ( X < 42000 ) = 1 - 0.9772 = 0.0228
:
**********************************************
The probability is greater with Company C that
you will get an offer of $42000 or greater
**********************************************
: